OpenStudy (anonymous):

2x - y - 4 = 0 3x + y - 9 = 0 What is the solution set of the given system?

OpenStudy (campbell_st):

just add the equations together, this will eliminate y and allow you to solve for x. when you have a value for x, substitute into either equation to find y

OpenStudy (anonymous):

how would i add them together?

OpenStudy (uri):

2x - y - 4 = 0 3x + y - 9 = 0 ======== 5x-13 x=13/5

OpenStudy (uri):

Now you have x you have to solve for y

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

would it be 6/5?

OpenStudy (uri):

Put the value of x in equation no2 3x + y - 9 = 0 3(13/5)+y-9=0 and yes...y =6/5

OpenStudy (anonymous):

oh ok thank you so much!

OpenStudy (uri):

You're welcome :)